# AMix-1: A Pathway to Test-Time Scalable Protein Foundation Model

### v1

@@ Abstract
We introduce AMix-1, a powerful protein foundation model built on Bayesian Flow Networks and empowered by a systematic training methodology, encompassing pretraining scaling laws, emergent capability analysis, in-context learning mechanism, and test-time scaling algorithm. To guarantee robust scalability, we establish a predictive scaling law and reveal the progressive emergence of structural understanding via loss perspective, culminating in a strong 1.7-billion model. Building on this foundation, we devise a multiple sequence alignment (MSA)-based in-context learning strategy to unify protein design into a general framework, where AMix-1 recognizes deep evolutionary signals among MSAs and consistently generates structurally and functionally coherent proteins. This framework enables the successful design of a dramatically improved AmeR variant with an up to 50×50$\times$50 × activity increase over its wild type. Pushing the boundaries of protein engineering, we further empower AMix-1 with an evolutionary test-time scaling algorithm for in silico directed evolution that delivers substantial, scalable performance gains as verification budgets are intensified, laying the groundwork for next-generation lab-in-the-loop protein design.
@@ 1 Introduction
In this study, we propose AMix-1, a pathway methodology for scalable protein design (Figure 1). We first characterize a predictable scaling law for Bayesian Flow Networks [30] and reveal that structural understanding emerges naturally as training progresses, with which we pretrained a powerful 1.7B AMix-1. Then we leverage its in-context learning capability with multiple sequence alignments to unify structure- and function-guided protein design into a general framework. With AMix-1, we successfuly design a novel AmeR variant validated through wet-lab experiments, achieving a 50×50$\times$50 × activity increase over the wild type. To push the limits further, we fledge AMix-1 with an evolutionary test-time scaling (TTS) algorithm for in silico directed evolution, allowing plug-and-play metrics and delivering strong performance gains with increasing verification budgets.
@@
  • We design a high-activity AmeR variant with a 50×50$\times$50 × improvement over the wild type, showcasing AMix-1’s real-world impact on functional protein engineering.
@@ 2 Protein Modeling with Bayesian Flow Network
In practice, we use the continuous categorical values of output distribution ${\bm{\phi}}(\boldsymbol{{\bm{\theta}}}_{i-1})$ without sampling to directly update ${\bm{\theta}}$ bypassing the sampling of noisy data needed for Bayesian update, which is similar to the noise reduced sampling method used in MolCRAFT [31]. The effective sampling method is described in appendix Algorithm 2. We also provide Figure 2 to help the readers visually grasp the key concepts of BFNs.
@@ 3.1 Pretraining
The scaling law exploration in the pre-training stage is conducted using the UniRef50 dataset, derived from UniProtKB [2] and selected UniParc sequences through iterative clustering (UniProtKB$+$UniParc $\rightarrow$ UniRef100 $\rightarrow$ UniRef90 $\rightarrow$ UniRef50) [32, 33]. This process ensures that UniRef50’s representative sequences are high-quality, non-redundant, and diverse, providing broad coverage of the protein sequence space for protein language models. Specifically, we utilized the pre-processed UniRef50 dataset provided by EvoDiff [34], comprising $41,546,293$ sequences for training and $82,929$ sequences for validation. Sequences longer than $1,024$ residues were randomly cropped to $1,024$ to reduce computational costs and generate diverse subsequences.
@@
Model Size $8$M $35$M $150$M $350$M $650$M $1.7$B
Number of layers $6$ $12$ $30$ $33$ $33$ $48$
@@
Attention heads $20$ $20$ $20$ $20$ $20$ $40$
AMix-1 model series are built upon the encoder-only Transformer architecture [35]. To encode positional information while preserving sequence invariance in attention patterns, we incorporate Rotary Position Embeddings (RoPE) [36], which introduce relative position dependencies through a rotation-based mechanism applied to query and key vectors. To explore the impact of model capacity on performance, we scale the architecture by systematically modifying the number of encoder layers, the number of attention heads, and the dimensionality of the hidden representations. These architectural configurations and corresponding parameter counts are summarized in Table 1.
@@
All models were trained using the Adamw [37] optimizer with a learning rate of $4\times 10^{-4}$, $\alpha$ values of $(0.9,0.98)$, and a weight decay of $0.01$. A polynomial learning rate schedule was adopted, spanning a total of $1$ million training steps with a warmup phase of $2,000$ steps. During warmup, the learning rate increased linearly from $1\times 10^{-7}$ to $4\times 10^{-4}$, followed by a polynomial decay to $4\times 10^{-5}$ with a power factor of $1$. Training was performed using mixed-precision (BF16) and a distributed data parallel (DDP) setup. To mitigate exploding gradients, gradient norms were clipped to a maximum of $1.0$. Early stopping was applied based on validation loss, with a patience threshold of $1,000$ steps. For AMix-1-1.7B and AMix-1-650M, training was conducted with a batch size of $4$ million across 16 nodes, each equipped with $8$ A800800800800 GPUs. Smaller models were trained with a batch size of $1$ million.
@@ 3.2 Predictive Scaling Behavior
Second, predictable loss modeling is closely linked to the phenomenon of emergence (see Section 4), in which specific downstream capabilities arise abruptly once the training loss falls below a critical threshold. Consequently, forecasting the loss trajectory not only enhances computational efficiency but also offers a principled approach to anticipating the onset of emergent behaviors.
@@
As for the metric, although we employ KL divergence (i.e., Equation 4) as the training objective, we report the cross-entropy between the output distribution $p_{\text{out}}$ and the ground-truth sequence, $\mathcal{L}=-\log p_{\text{out}}$, to evaluate scaling behavior. This choice is motivated by two factors. First, from an information-theoretic perspective, the cross-entropy measures the inefficiency of encoding the ground-truth sequence using the model’s predicted distribution. Since the ground truth is inherently one-hot, minimizing cross-entropy directly reflects the model’s ability to recover precise, token-level information. Second, prior works on scaling laws [25, 26, 42] have consistently adopted cross-entropy, allowing for direct and meaningful comparisons.
@@
where $E_{0}$ denotes the irreducible loss, representing the asymptotic lower bound that cannot be reduced through additional compute. As shown in Figure 3, our AMix-1 models exhibit clear scaling behavior under two noise levels $\alpha=0.16$ and $0.32$: the validated data points (red stars) align closely with the extrapolated power-law curves, consistent with the scaling laws observed in large language models [42, 19]. This consistency enables reliable performance forecasting and informs efficient compute allocation for larger training regimes.
In addition to moderate noise levels ($\alpha=0.16$ and $0.32$), we evaluate $\mathcal{L}$ under extreme noise conditions ($\alpha=0.08$ and $0.64$), as shown in Figure 13. Under these settings, the BFN models fail to capture the training objective defined in Equation 4, leading to erratic and unpredictable cross-entropy values. What’s more, although both $\alpha=0.16$ and $\alpha=0.32$ yield similar scaling trends, the subtle differences in their exponents indicate that careful scheduling of $\alpha$ can significantly influence convergence behavior and generalization. Since $\alpha$ directly modulates the injected noise, we point out that an appropriate schedule of $\alpha$ when training may be essential for stabilizing training and achieving optimal performance in BFN-based models.
@@
$\alpha$ $E$ $A$ $B$ $n$ $d$ Fitting Valid
@@ 3.3 Optimal Estimation under Fixed FLOPs
As discussed in Section 3.2, we investigate the parametric scaling law of AMix-1 models across different noise levels $\alpha$, with detailed results summarized in Table 2. Additional fitting results for other noise settings are provided in Section 11.3. To ensure robustness and assess generalization, we split the data by training tokens, using the former $30\%$ for fitting and the latter $30\%$ for validation. Following Hoffmann et al. [26], we minimize the Huber loss [43] between the predicted and observed log loss using the L-BFGS algorithm [44] to estimate Equation 7. We quantify the goodness of fit using two metrics: the Fitting Mean Relative Error (MRE) and the Validation MRE, defined in Section 11.2. These metrics reflect the relative prediction accuracy within and beyond the fitting region, respectively.
@@
For each noise level $\alpha$ under which BFN models exhibit scalable behavior, we analyze the compute-optimal Pareto frontier to identify the most efficient combination of model size and training token count that minimizes validation loss under a fixed computational budget. This frontier delineates the best attainable trade-off between parameter count and data usage, offering a principled framework for resource allocation across varying noise regimes. Under a given noise level $\alpha$, our objective is to characterize the compute-optimal frontier that relates the minimum achievable cross-entropy $\hat{\mathcal{L}}$ to the training FLOPs $F$, similar to Equation 6:
^$\hat{\mathcal{L}}=E_{0}+e^{a}\cdot F^{-b}.$ (8)
To accurately estimate the scaling parameters $a$ and $b$, we analyze the cross-entropy loss as a function of FLOPs for six AMix-1 model variants ranging from 8M to 1.7B parameters. In Figure 4, we present two compute-optimal frontiers for $\alpha=0.16$ and $\alpha=0.32$. Each curve corresponds to a fixed model size, with dots denoting empirical measurements under different training regimes and solid lines representing smoothed scaling trends. Larger models such as AMix-1-650M and AMix-1-1.7B display steeper loss reductions, suggesting greater efficiency in leveraging additional compute. In contrast, smaller models, such as AMix-1-8M and AMix-1-35M, saturate earlier, indicating a limited capacity to benefit from extended training. The dashed black line shows the global Pareto frontier, fit across all configurations. Notably, AMix-1-650M and AMix-1-1.7B closely align with this frontier, demonstrating near-optimal compute efficiency. Conversely, the pronounced deviation of smaller models from the frontier highlights their sub-optimality in the compute–performance trade-off.
@@ 4 Emergent Ability
Building on the predictive scaling law established in Section 3, we further investigate the less predictive aspects of AMix-1’s performance in protein generation. Our analysis of emergent behavior adopts a loss-centered perspective, following Du et al. [46]. We adopt the predictive cross-entropy loss in Section 3 as an anchor to empirically map training loss to protein generation performance. This empirical mapping provides a practical framework for interpreting how emergent capabilities arise during training, offering insight into the relationship between optimization dynamics and functional outcomes in protein modeling.
@@ 4.1 Experimental Setup
In protein sequence learning, a central question is how structural information emerges as sequences are progressively learned, an inquiry grounded in the widely accepted “sequence-structure-function” paradigm [47]. To investigate this emergent process, we focus on three key capabilities: sequence consistency, reflecting sequence-level recovery from corrupted sequence distribution; foldability, representing the transition from sequence understanding to structural feasibility; and structural consistency, indicating the model’s ability to preserve structural features. Notably, we demonstrate that structural consistency plays a critical role, as it underpins the emergence of few-shot learning capabilities discussed in Section 5.
@@
For the above metrics, we take Automated Model Evaluation (CAMEO) [49] as our evaluation benchmark. It is a continuously updated benchmark dataset that provides newly released, high-quality protein structures from the Protein Data Bank (PDB) for real-time assessment of protein structure prediction methods. To assess the scalability of our models, we randomly selected $20$ proteins from CAMEO. Details of these sequences are provided in Section 10.
In this emergent ability analysis, AMix-1 adopts a setup consistent with the scaling law experiments described in  Section 3. We do not apply additional decoding strategies, unlike previous foundation models [50, 51], in order to stabilize the emergent ability analysis and avoid introducing confounding effects. We structure the analysis along the axis of corruption noise, denoted as $\alpha$. Specifically, we investigate four representative values: $\{0.08,0.16,0.48,0.64\}$, where $\alpha=0.08$ and $\alpha=0.64$ correspond to extremely high and low noise levels, respectively, providing a comprehensive view of emergent behavior across the noise spectrum. These same $\alpha$ values are also used as the initial noise levels in AMix-1’s generation process. Our goal is to align the cross-entropy dynamics at these specific noise levels with the corresponding generation behavior, enabling a direct mapping between training metrics and generation performance.
@@ 4.2 Findings
We conduct emergent ability analysis using sequence consistency, foldability, and structural consistency as evaluation axes, and use the starting noise levels $\alpha\in\{0.08,0.16,0.48,0.64\}$ as the noise dimension. The results are presented in Figure 5, from which we draw several key findings summarized below.
@@ 5 In-Context Learning
In-context learning (ICL) emerges as a central capability that enables large language models learns from a few contextual exemplars to perform unlimited tasks without task-specific fine-tuning [17, 52]. This intriguing ability motivated us to develop a training-free strategy to unify diverse protein design tasks by leveraging the in-context learning ability of AMix-1, which also serves as the foundation of our test-time scaling algorithm described in Section 6. In this section, we propose a novel approach that conditions AMix-1 on evolutionary profiles as prompts, enabling the generation of proteins that preserve both structural integrity and functional relevance. We validated the effectiveness and generalizability of this approach through extensive in-silico evaluations in structural and functional consistency. Wet-lab experiments further confirmed its potential, successfully engineering the AmeR transcriptional repressor with 50-fold increase in activity over the wild type.

### v3

@@ Abstract
We introduce AMix-1, a powerful protein foundation model built on Bayesian Flow Networks and empowered by a systematic training methodology, encompassing pretraining scaling laws, emergent capability analysis, in-context learning mechanism, and test-time scaling algorithm. To guarantee robust scalability, we establish a predictive scaling law and reveal the progressive emergence of structural understanding via loss perspective, culminating in a strong 1.7-billion model. Building on this foundation, we devise a multiple sequence alignment (MSA)-based in-context learning strategy to unify protein design into a general framework, where AMix-1 recognizes deep evolutionary signals among MSAs and consistently generates structurally and functionally coherent proteins. This framework enables the successful design of a dramatically improved AmeR variant with an up to $50\times$ activity increase over its wild type. Pushing the boundaries of protein engineering, we further empower AMix-1 with an evolutionary test-time scaling algorithm for in silico directed evolution that delivers substantial, scalable performance gains as verification budgets are intensified, laying the groundwork for next-generation lab-in-the-loop protein design.
@@ 1 Introduction
In this study, we propose AMix-1, a pathway methodology for scalable protein design (Figure˜1). We first characterize a predictable scaling law for Bayesian Flow Networks [30] and reveal that structural understanding emerges naturally as training progresses, with which we pretrained a powerful 1.7B AMix-1. Then we leverage its in-context learning capability with multiple sequence alignments to unify structure- and function-guided protein design into a general framework. With AMix-1, we successfuly design a novel AmeR variant validated through wet-lab experiments, achieving a $50\times$ activity increase over the wild type. To push the limits further, we fledge AMix-1 with an evolutionary test-time scaling (TTS) algorithm for in silico directed evolution, allowing plug-and-play metrics and delivering strong performance gains with increasing verification budgets.
@@
  • We design a high-activity AmeR variant with a $50\times$ improvement over the wild type, showcasing AMix-1’s real-world impact on functional protein engineering.
@@ 2 Protein Modeling with Bayesian Flow Network
In practice, we use the continuous categorical values of output distribution ${\bm{\phi}}(\boldsymbol{{\bm{\theta}}}_{i-1})$ without sampling to directly update $\bm{\theta}$} bypassing the sampling of noisy data needed for Bayesian update, which is similar to the noise reduced sampling method used in MolCRAFT [31]. The effective sampling method is described in appendix Algorithm 2. We also provide Figure˜2 to help the readers visually grasp the key concepts of BFNs.
@@ 3.1 Pretraining
The scaling law exploration in the pre-training stage is conducted using the UniRef50 dataset, derived from UniProtKB [2] and selected UniParc sequences through iterative clustering (UniProtKB++UniParc $\rightarrow$ UniRef100 $\rightarrow$ UniRef90 $\rightarrow$ UniRef50) [32, 33]. This process ensures that UniRef50’s representative sequences are high-quality, non-redundant, and diverse, providing broad coverage of the protein sequence space for protein language models. Specifically, we utilized the pre-processed UniRef50 dataset provided by EvoDiff [34], comprising $41,546,293$ sequences for training and $82,929$ sequences for validation. Sequences longer than $1,024$ residues were randomly cropped to $1,024$ to reduce computational costs and generate diverse subsequences.
@@
Model Size 88M 3535M 150150M 350350M 650650M 1.71.7B
Number of layers 66 $12$ $30$ $33$ $33$ $48$
@@
Attention heads 2020 2020 2020 2020 2020 $40$
AMix-1 model series are built upon the encoder-only Transformer architecture [35]. To encode positional information while preserving sequence invariance in attention patterns, we incorporate Rotary Position Embeddings (RoPE) [36], which introduce relative position dependencies through a rotation-based mechanism applied to query and key vectors. To explore the impact of model capacity on performance, we scale the architecture by systematically modifying the number of encoder layers, the number of attention heads, and the dimensionality of the hidden representations. These architectural configurations and corresponding parameter counts are summarized in Table˜1.
@@
All models were trained using the Adamw [37] optimizer with a learning rate of $4\times 10^{-4}$, $\alpha$ values of $(0.9,0.98)$, and a weight decay of $0.01$. A polynomial learning rate schedule was adopted, spanning a total of $1$ million training steps with a warmup phase of $2,000$ steps. During warmup, the learning rate increased linearly from $1\times 10^{-7}$ to $4\times 10^{-4}$, followed by a polynomial decay to $4\times 10^{-5}$ with a power factor of $1$. Training was performed using mixed-precision (BF16) and a distributed data parallel (DDP) setup. To mitigate exploding gradients, gradient norms were clipped to a maximum of $1.0$. Early stopping was applied based on validation loss, with a patience threshold of $1,000$ steps. For AMix-1-1.7B and AMix-1-650M, training was conducted with a batch size of $4$ million across 16 nodes, each equipped with 88 A800800 GPUs. Smaller models were trained with a batch size of $1$ million.
@@ 3.2 Predictive Scaling Behavior
Second, predictable loss modeling is closely linked to the phenomenon of emergence (see Section˜4), in which specific downstream capabilities arise abruptly once the training loss falls below a critical threshold. Consequently, forecasting the loss trajectory not only enhances computational efficiency but also offers a principled approach to anticipating the onset of emergent behaviors.
@@
As for the metric, although we employ KL divergence (i.e., Equation˜4) as the training objective, we report the cross-entropy between the output distribution $p_{\text{out}}$ and the ground-truth sequence, $\mathcal{L}=-\log p_{\text{out}}$, to evaluate scaling behavior. This choice is motivated by two factors. First, from an information-theoretic perspective, the cross-entropy measures the inefficiency of encoding the ground-truth sequence using the model’s predicted distribution. Since the ground truth is inherently one-hot, minimizing cross-entropy directly reflects the model’s ability to recover precise, token-level information. Second, prior works on scaling laws [25, 26, 42] have consistently adopted cross-entropy, allowing for direct and meaningful comparisons.
@@
where $E_{0}$ denotes the irreducible loss, representing the asymptotic lower bound that cannot be reduced through additional compute. As shown in Figure˜3, our AMix-1 models exhibit clear scaling behavior under two noise levels $\alpha=0.16$ and $0.32$: the validated data points (red stars) align closely with the extrapolated power-law curves, consistent with the scaling laws observed in large language models [42, 19]. This consistency enables reliable performance forecasting and informs efficient compute allocation for larger training regimes.
In addition to moderate noise levels ($\alpha=0.16$ and $0.32$), we evaluate $\mathcal{L}$ under extreme noise conditions ($\alpha=0.08$ and $0.64$), as shown in Figure˜13. Under these settings, the BFN models fail to capture the training objective defined in Equation˜4, leading to erratic and unpredictable cross-entropy values. What’s more, although both $\alpha=0.16$ and $\alpha=0.32$ yield similar scaling trends, the subtle differences in their exponents indicate that careful scheduling of $\alpha$ can significantly influence convergence behavior and generalization. Since $\alpha$ directly modulates the injected noise, we point out that an appropriate schedule of $\alpha$ when training may be essential for stabilizing training and achieving optimal performance in BFN-based models.
@@
$\alpha$ $E$ AA $B$ $n$ $d$ Fitting Valid
@@ 3.3 Optimal Estimation under Fixed FLOPs
As discussed in Section˜3.2, we investigate the parametric scaling law of AMix-1 models across different noise levels $\alpha$, with detailed results summarized in Table˜2. Additional fitting results for other noise settings are provided in Section˜11.3. To ensure robustness and assess generalization, we split the data by training tokens, using the former $30\%$ for fitting and the latter $30\%$ for validation. Following Hoffmann et al. [26], we minimize the Huber loss [43] between the predicted and observed log loss using the L-BFGS algorithm [44] to estimate Equation˜7. We quantify the goodness of fit using two metrics: the Fitting Mean Relative Error (MRE) and the Validation MRE, defined in Section˜11.2. These metrics reflect the relative prediction accuracy within and beyond the fitting region, respectively.
@@
For each noise level $\alpha$ under which BFN models exhibit scalable behavior, we analyze the compute-optimal Pareto frontier to identify the most efficient combination of model size and training token count that minimizes validation loss under a fixed computational budget. This frontier delineates the best attainable trade-off between parameter count and data usage, offering a principled framework for resource allocation across varying noise regimes. Under a given noise level $\alpha$, our objective is to characterize the compute-optimal frontier that relates the minimum achievable cross-entropy $\hat{\mathcal{L}}$ to the training FLOPs $F$, similar to Equation˜6:
$\hat{\mathcal{L}}=E_{0}+e^{a}\cdot F^{-b}.$ (8)
To accurately estimate the scaling parameters $a$ and $b$, we analyze the cross-entropy loss as a function of FLOPs for six AMix-1 model variants ranging from 8M to 1.7B parameters. In Figure˜4, we present two compute-optimal frontiers for $\alpha=0.16$ and $\alpha=0.32$. Each curve corresponds to a fixed model size, with dots denoting empirical measurements under different training regimes and solid lines representing smoothed scaling trends. Larger models such as AMix-1-650M and AMix-1-1.7B display steeper loss reductions, suggesting greater efficiency in leveraging additional compute. In contrast, smaller models, such as AMix-1-8M and AMix-1-35M, saturate earlier, indicating a limited capacity to benefit from extended training. The dashed black line shows the global Pareto frontier, fit across all configurations. Notably, AMix-1-650M and AMix-1-1.7B closely align with this frontier, demonstrating near-optimal compute efficiency. Conversely, the pronounced deviation of smaller models from the frontier highlights their sub-optimality in the compute–performance trade-off.
@@ 4 Emergent Ability
Building on the predictive scaling law established in Section˜3, we further investigate the less predictive aspects of AMix-1’s performance in protein generation. Our analysis of emergent behavior adopts a loss-centered perspective, following Du et al. [46]. We adopt the predictive cross-entropy loss in Section˜3 as an anchor to empirically map training loss to protein generation performance. This empirical mapping provides a practical framework for interpreting how emergent capabilities arise during training, offering insight into the relationship between optimization dynamics and functional outcomes in protein modeling.
@@ 4.1 Experimental Setup
In protein sequence learning, a central question is how structural information emerges as sequences are progressively learned, an inquiry grounded in the widely accepted “sequence-structure-function” paradigm [47]. To investigate this emergent process, we focus on three key capabilities: sequence consistency, reflecting sequence-level recovery from corrupted sequence distribution; foldability, representing the transition from sequence understanding to structural feasibility; and structural consistency, indicating the model’s ability to preserve structural features. Notably, we demonstrate that structural consistency plays a critical role, as it underpins the emergence of few-shot learning capabilities discussed in Section˜5.
@@
For the above metrics, we take Automated Model Evaluation (CAMEO) [49] as our evaluation benchmark. It is a continuously updated benchmark dataset that provides newly released, high-quality protein structures from the Protein Data Bank (PDB) for real-time assessment of protein structure prediction methods. To assess the scalability of our models, we randomly selected $20$ proteins from CAMEO. Details of these sequences are provided in Section˜10.
In this emergent ability analysis, AMix-1 adopts a setup consistent with the scaling law experiments described in  Section˜3. We do not apply additional decoding strategies, unlike previous foundation models [50, 51], in order to stabilize the emergent ability analysis and avoid introducing confounding effects. We structure the analysis along the axis of corruption noise, denoted as $\alpha$. Specifically, we investigate four representative values: $\{0.08,0.16,0.48,0.64\}$, where $\alpha=0.08$ and $\alpha=0.64$ correspond to extremely high and low noise levels, respectively, providing a comprehensive view of emergent behavior across the noise spectrum. These same $\alpha$ values are also used as the initial noise levels in AMix-1’s generation process. Our goal is to align the cross-entropy dynamics at these specific noise levels with the corresponding generation behavior, enabling a direct mapping between training metrics and generation performance.
@@ 4.2 Findings
We conduct emergent ability analysis using sequence consistency, foldability, and structural consistency as evaluation axes, and use the starting noise levels $\alpha\in\{0.08,0.16,0.48,0.64\}$ as the noise dimension. The results are presented in Figure˜5, from which we draw several key findings summarized below.
@@ 5 In-Context Learning
In-context learning (ICL) emerges as a central capability that enables large language models learns from a few contextual exemplars to perform unlimited tasks without task-specific fine-tuning [17, 52]. This intriguing ability motivated us to develop a training-free strategy to unify diverse protein design tasks by leveraging the in-context learning ability of AMix-1, which also serves as the foundation of our test-time scaling algorithm described in Section˜6. In this section, we propose a novel approach that conditions AMix-1 on evolutionary profiles as prompts, enabling the generation of proteins that preserve both structural integrity and functional relevance. We validated the effectiveness and generalizability of this approach through extensive in-silico evaluations in structural and functional consistency. Wet-lab experiments further confirmed its potential, successfully engineering the AmeR transcriptional repressor with 50-fold increase in activity over the wild type.